Air Pollution the amount of mitrogen dioxide A in parts per million (PPM) that presetn in the air in the city of homer on a certain day in june is modeled by the function a(t)=-2t^2+32t+12
Where t is the number of hours after 6:00 am use this function to find the time at which the nitrogen dioxide level was at its maximum.

Answer by nerdybill(7384)

Air Pollution the amount of mitrogen dioxide A in parts per million (PPM) that presetn in the air in the city of homer on a certain day in june is modeled by the function a(t)=-2t^2+32t+12
Where t is the number of hours after 6:00 am use this function to find the time at which the nitrogen dioxide level was at its maximum.
.
Because the coefficient associated with the t^2 term is negative, we KNOW that it is a parabola that opens downwards. So, the vertex is the max.
.
t-value at max:
t = -b/(2a)
t = -32/(2(-2))
t = -32/(-4)
t = 8 hours
.
6:00 am + 8 hours = 2 pm